OpenStudy (tt1618):

Adiana has a rectangular garden that measures 9 m wide and 13 m long. She wants to increase the area to 192 m^2 by increasing the width and the height by the same length. What will be the new length of the garden? A. 16 m B. 15 m C. 14 m D. 13 m

OpenStudy (dayakar):

length of first rectangle = 13m , width = 9m area = 13 *9 =117 m^2-----(1) increase in area = 192m^2----(2) area of the new rectangle = 117 + 192 = 364 m^2----(3) i l = 13 , width = w area = 364 m^2 l*w =364 13 * w= 364 find the width ? and find the difference between first width and second width

OpenStudy (dayakar):

do u understand !

OpenStudy (tt1618):

I think so, so are you saying the answer is 13 m?

OpenStudy (matthew120):

yea its 13

OpenStudy (tt1618):

Thanks

OpenStudy (matthew120):

np

OpenStudy (tt1618):

It was 16

OpenStudy (dayakar):

w = 364/13 new width = 28 length is 13 , it is same according to the problem

OpenStudy (retireed):

You misread the question. The NEW area was 192, so X the increase in both length and width is 3. Solving for the root that makes sense. The new length is 13 + 3 = 16